Half-Life

The half-life of carbon-14 is 5715 years. If an animal bone contains 20.0 grams of carbon-14 when the animal dies, how many grams of carbon-14 will be left in 17145years (3 half-lives)?
Answer #1

After 5715 years (1 half life)

Mass remaining = 20.0/2 = 10.0 g

After 11430 years (2 half lives)

Mass remaining = 10.0/2 = 5.0 g

After 17145 years (3 half lives)

Mass remaining = 5.0/2 = 2.5 g

answered by: glloyd